Two trainers(Anna and Dan) place an injured athlete on a stretcher. They carry Tristan from the field. Tristan weighs 1432 N. The stretcher is 2.25 M long and weighs 25N. That puts Tristan at the center of mass 1.0 m from Dan’s end of the stretcher.

how much force did anna and dan have to exert to keep the man leveled?

Please use the GUESS method.

Thanks.
Answer #1

I don't know what the GUESS method is...

I will use my own method ................ Remember, the method you use in class may not be familiar to all.....

------------------------------------------------------------------------

this is what I made based on question description

dan Anna 25 N Fd 1432 N 1.25 m

Fd is the force applied by Dan and Fa is force applied by Anna

take torque about right end, we have

Fd * 2.25 = 1432 * 1.25 + 25 * 1.125

Fd = 808.1 N

-----------------

now , equate forces in vertical direction

Fd + Fa = 1432 + 25

Fa = 1432 + 25 - Fd

Fa = 649 N